Use the following information, evaluate the 1998 NHL Lockout

1) The Levitt report shows NHL revenues at 2.094 billion in 1998, throughout the lockout, the owners claimed that players were receiving 75% of the revenue. When the new CBA settled, players' revenue was reduced to 54%. The new contract lasted 6 years long. Using the information above, calculate the owner's gain from this concession.

2) Assume the league's profits in a season was $335 million. Assume costs of fan irritation from the lockout is equal to owners' next best alternative, calculate the owners' break-even probability. Explain your results.
Answer #1

ANSWER

As per the given question,

NHL revenues were 2.094 billion.

Players were receiving 75% of the revenue, therefore amount = .75*2.094 billions= 1.5705 billions

Players revenue at 54% = .54*2.094 = 1.13076 Billions

Therefore, owners's gain from the concession = 1.5705 - 1.13076 = .43974 billions= 439.74 millions.